OpenStudy (ivancsc1996):

Can somebody help me writing \(\int \frac{x-\tan^{-1}(x)}{x^{3}}\textrm{d}x\) as a power series?

OpenStudy (anonymous):

For starters, let's integrate: \[\int\frac{x-\arctan x}{x^3}\,dx=\int\frac{dx}{x^2}-\int\frac{\arctan x}{x^3}\,dx\] For the second integral, integrate by parts with \[\begin{matrix} u=\arctan x&&&dv=\dfrac{dx}{x^3}\\ du=\dfrac{dx}{1+x^2}&&&v=-\dfrac{1}{2x^2} \end{matrix}\] So you have \[\int\frac{dx}{x^2}-\left(-\frac{\arctan x}{2x^2}+\frac{1}{2}\int\frac{dx}{x^2(1+x^2)}\right)\] \[-\frac{1}{x}+\frac{\arctan x}{2x^2}-\frac{1}{2}\int\frac{dx}{x^2(1+x^2)}\] The remaining integral can be dealt with via partial fraction decomposition - I'll leave the work to you in verifying that the above is equivalent to \[-\frac{1}{2x}+\frac{\arctan x}{2x^2}+\frac{1}{2}\arctan x\] Now let's worry about the power series representation.

OpenStudy (anonymous):

Let's consider the derivative of \(\arctan x\): \[\frac{d}{dx}\arctan x=\frac{1}{1+x^2}=\frac{1}{1-(-x^2)}\] For \(|-x^2|=|x|^2<1\), or simply \(|x|<1\), the above has the geometric series representation, \[\frac{d}{dx}\arctan x=\sum_{k=0}^\infty (-x^2)^k=\sum_{k=0}^\infty (-1)^kx^{2k}\] Take the antiderivative will give us the power series for \(\arctan x\): \[\begin{align*}\int\frac{d}{dx}\arctan x\,dx&=\int\sum_{k=0}^\infty (-1)^kx^{2k}\,dx\\\\ \arctan x&=\int\sum_{k=0}^\infty (-1)^kx^{2k}\,dx\\\\ &=\sum_{k=0}^\infty (-1)^k\int x^{2k}\,dx\\\\ &=\sum_{k=0}^\infty (-1)^k\left(\frac{x^{2k+1}}{2k+1}+C\right)\end{align*}\] We know that for \(x=0\), we have \(\arctan x=0\), so \(C\) must also be \(0\). (Alternatively, we can reason that if \(C\) were non-zero, the series would not converge, so it must be zero.)

OpenStudy (anonymous):

So you can finally replace \(\arctan x\) in original integral result and you'll have your very own power series for \(\displaystyle\int\frac{x-\arctan x}{x^3}\,dx\).